Calculating free cash flows ) At present, Solartech Skateboards is considering expanding its product line to include gas-powered skateboards; however, it is questionable how well they will be received by skateboarders. Although you feel there is a 60 percent chance you will sell 8,000 of these per year for 10 years (after which time this project is expected to shut down because solar-powered skateboards will become more popular), you also recognize that there is a 20 percent chance that you will only sell 4,000 and also a 20 percent chance you will sell 15,000. The gas skateboards would sell for $ 100 each and have a variable cost of $30 each. Regardless of how many you sell, the annual fixed costs associated with production would be $120,000. In addition, there would be an initial expenditure of $1,100,000 associated with the purchase of new production equipment. It is assumed that this initial expenditure will be depreciated using the simplified straight-line method down to zero over 10 years. Because of the number of stores that will need inventory, the working capital requirements are the same regardless of the level of sales. This project will require a one-time initial investment of $60,000 in net working capital, and that working-capital investment will be recovered when the project is shut down. Finally, assume that the firm's marginal tax rate is 32 percent.

***Please Show All Work (in the most simplistic form clearly illustrated)***

a. What is the initial outlay associated with this project? (Round to the nearest dollar.)

b. If your sales forecast is 8,000 skateboards per year, what are the annual free cash flows associated with this project for years 1 through 9 (note that the cash flows for years 1 through 9 are equal)? (Round to the nearest dollar.)

If your sales forecast is 4,000 skateboards per year, what are the annual free cash flows associated with this project for years 1 through 9 (note that the cash flows for years 1 through 9 are equal)? (Round to the nearest dollar.)

If your sales forecast is 15,000 skateboards per year, what are the annual free cash flows associated with this project for years 1 through 9 (note that the cash flows for years 1 through 9 are equal)? (Round to the nearest dollar.)

What are the expected annual free cash flows for years 1 through 9 (note that the cash flows for years 1 through 9 are equal)?(Round to the nearest dollar.)

c. If your sales forecast is 8,000 skateboards per year, what is the terminal cash flow in year 10 (that is, what is the free cash flow in year 10 plus any additional cash flows associated with the termination of the project)? (Round to the nearest dollar.)

If your sales forecast is 4,000 skateboards per year, what is the terminal cash flow in year 10 (that is, what is the free cash flow in year 10 plus any additional cash flows associated with the termination of the project)? (Round to the nearest dollar.)

If your sales forecast is 15,000 skateboards per year, what is the terminal cash flow in year 10 (that is, what is the free cash flow in year 10 plus any additional cash flows associated with the termination of the project)? (Round to the nearest dollar.)

What is the expected terminal cash flow in year 10? (Round to the nearest dollar.)

d. Using the expected free cash flows, what is the project's NPV given a required rate of return of 11 percent? (Round to the nearest dollar.)

What would the project's NPV be if they sold 8,000 skateboards? (Round to the nearest dollar.)
Answer #1

Answer:

(a) The initial outlay includes the cost of new machine as well as the investment in working captal.

$1,100,000 + $60,000 = $1,160,000

(b) Earning before interst and tax

EBIT = Sales * (Actual Price - Variable Price) - Fixed Cost - Depreciation

EBIT = 8000 * (100 - 30) - 120000 - 110000

EBIT = 330,000

After taxes, the net income is = 330,000 (1 - 0.32) = 224,400

EBIT = 4000 * (100 - 30) - 120000 - 110000

EBIT = 50,000

After taxes, the net income is = 50,000 (1 - 0.32) = 34,000

EBIT = 15000 * (100 - 30) - 120000 - 110000

EBIT = 820,000

After taxes, the net income is = 330,000 (1 - 0.32) = 557,600

Annual free cash flows for years 1 through 9 is

(0.60) * 224,400 + (0.20) * 34,000 + (0.20) * 557,600

= 134,640 + 6,800 + 111,520 = 252,960

Adding back the depreciation to find the annual free cash flow:

252960 + 110000 = $362960

(c) Terminal cash flow in 10 years will be annual year cash flow plus initial working capital

362960 + 60000 = $422960

(d) NPV = -Initial cost + annual cash flow/ (1 + r)1 + annual cash flow/ (1 + r)2 +.....+ annual cash flow/ (1 + r)9 + Terminal cah flow/ (1 + r)10

NPV = -1160000 + 362960/ 1.111 + 362960/ 1.112 +....+ 362960/ 1.119 + 422960/ 1.1110

NPV = -1160000 + 326990.99 + 294586.48 + 265321.64 + 239104.08 + 215406.53 + 194096.26 + 174836.22 + 157466.38 + 141892.10 + 148982.04

NPV = 998682.72 = 998683